Determine the total capacitance if a 30 pF capacitor is connected in series with 0.03µF  capacitor. 


1
Expert's answer
2022-02-28T10:04:54-0500

Solution;

For capacitors in series;

"\frac{1}{C_T}=\frac{1}{C_1}+\frac{1}{C_2}"

"\frac{1}{C_T}=\frac{1}{30×10^{-12}}+\frac{1}{0.03×10^{-6}}"

"C_T=2.997\u00d710^{-11}F"